Find the cylindrical coordinates of the points where the Cartesian coordinates are :


(3, 3,4)


1
Expert's answer
2022-01-24T08:25:16-0500

Solution

Relationship between Cartesian coordinates and Cylindrical coordinates is as follows:

"r^2=x^2+y^2"

"tan\theta=\frac yx"

"z=z"

Given

x=3

y=3

z=4

Hence

"r^2=3^2+3^2=18"

"r=\sqrt{18}=3\sqrt2"

"tan(\theta)=\frac{3}{3}=1"

"\theta=\frac{π}{4}"

"z=4"

The cylindrical coordinates are given by

"(3\sqrt2,\frac{π}{4},4)" .